OpenStudy (anonymous):

cos pi/4 cos pi/6=1/2(___pi/12+cos5pi/12)

OpenStudy (jdoe0001):

hint: \(\bf cos({\color{brown}{ \alpha}})cos({\color{blue}{ \beta}})=\cfrac{1}{2}[cos({\color{brown}{ \alpha}}-{\color{blue}{ \beta}})\quad +\quad cos({\color{brown}{ \alpha}}+{\color{blue}{ \beta}})]\)

OpenStudy (anonymous):

cos?

OpenStudy (lynfran):

\[\cos (\alpha)\cos(\beta)=\frac{ 1 }{ 2 }[\cos(\alpha-\beta)+\cos(\alpha+\beta)]\]\[\cos \frac{ \pi }{ 4 }\cos \frac{ \pi }{ 6 }=\frac{ 1 }{ 2 }[\cos(\frac{ \pi }{ 4 }-\frac{ \pi }{ 6 })+\cos(\frac{ \pi }{4 }+\frac{ \pi }{ 6 })]\]

OpenStudy (lynfran):

yes cos

OpenStudy (anonymous):

YAY! Thanks:)

OpenStudy (lynfran):

welcome